OpenStudy (anonymous):

Evaluate the following Integral by using the substitution method: (gonna type it up after this is posted)

OpenStudy (anonymous):

\[\int\limits 1/x^2[\cos^2(1/x)] dx\]

OpenStudy (experimentx):

let 1/x = u, -1/x^2 dx = du

OpenStudy (anonymous):

and then? :)

OpenStudy (experimentx):